OpenStudy (anonymous):

Find exact value without using calculator. cos10/sin80

hartnn (hartnn):

use the relation, \(\Large \cos (90-x) =\sin x\) plug in x =80

OpenStudy (anonymous):

cos10=0.98 sin80=0.95 so just divide them=1.03

OpenStudy (anonymous):

hartnn: Would you mind explaining a bit more? I got values that are not on the pie chart. Thanks. zaibali: Can't use a calculator. Teacher doesn't allow calculators on test.

OpenStudy (anonymous):

oh...sorrey...i get like you write use calculator...thats why i write it....!!!

hartnn (hartnn):

did you try plugging in x=80 in the formula i gave ?? what did u get when u put x=80?

OpenStudy (anonymous):

I got cos(10)=sin(80)

hartnn (hartnn):

correct so cos 10 = sin 80 then cos 10/sin 80 will be ... ?

OpenStudy (anonymous):

They have to be the same number but I don't know the cos 10 or sin 80

hartnn (hartnn):

.-. if a=b then a/b =1 if cos 10 = sin 80 then cos 10/sin 80 =...

hartnn (hartnn):

you don't even need the actual values of cos 10 or sin 80

OpenStudy (anonymous):

Wow I'm so bad at math. I'm sorry for the frustration lol. It wasn't a good idea for me to take trig in the quick paced summer course.

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

just practice few problems, then you will be able to solve it properly :) and welcome ^_^

OpenStudy (anonymous):

\[1-\cos ^{2}20-\cos^270\] sin^2 20 - cos^2 70 sin^2 20 - sin^2 (90-70) sin^2 20 - sin^2 20 =0?

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

Thanks I think I got the hang of it.

hartnn (hartnn):

ofcourse! trigonometry is easy :)

OpenStudy (anonymous):

It is but there are parts where I just get lost since I don't understand what is going on.

hartnn (hartnn):

that happens, nothing to worry, if you're stuck, just ask! thats why we are here :) to exchange knowledge :)

OpenStudy (anonymous):

I'm glad I found this website. Everyone is friendly and very helpful!

hartnn (hartnn):

we are also glad to have you here :) \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

Thank you and I'm glad to be here :)